A hospital needs to dilute a 60% boric acid solution to a 10% solution. If it needs 40 liters of the 10% solution, how much of the 60% solution and how much water should it use?
Answer by edjones(8007)

40 * 1/6 = 40/6 = 6 2/3 liters.